# eprint Nisho-2-2025

### v1

@@ Abstract
A dark matter axion with mass $m_{a}$ induces an oscillating electric field in a cylindrical sample placed under a magnetic field $B$ parallel to the cylinder axis. When the cylinder is made of a highly electrically conductive material, the induced oscillating current primarily dissipates the axion energy at the surface. In contrast, if the cylinder is composed of a material with low conductivity, e.g. $\sigma=10^{-3}\text{eV}$, the axion energy is dissipated mainly inside the bulk of the cylinder. Within the QCD axion model, the dissipated power $P$ is estimated as $P\simeq 2.8\times 10^{-28}\text{W}g_{\gamma}^{2}\left(L/100\text{cm}\right)\left(R/2\text{cm}\right)^{2}\left(m_{a}/10^{-4}\text{eV}\right)\left(B/10\text{T}\right)^{2}\left(10/\epsilon\right)\left(\rho_{a}/0.3\text{GeVcm}^{-3}\right)$, with $\sigma=10^{-3}$eV, radius $R$, length $L$, electric permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$. Using an LC circuit appropriately tuned to a quality factor $Q=10^{6}$, the signal-to-noise ratio at low temperature 100100mK and observation time 6060sec. is so large $\sim 10$ that the detection of dark matter axions is feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
@@
In this letter we propose a new method for axion detection using a cylindrical sample with low electric conductivity, $\sigma=10^{-3}\text{-}10^{-2}\text{eV}$ at low temperature 100100mK. Our target is the mass range $m_{a}=10^{-4}\mbox{-}10^{-3}$eV, which is difficult to detect with resonant cavity.
@@
In general, such currents are confined within the skin depth $\delta$. For highly conductive materials, e.g. $\sigma=10^{4}\text{eV}$, the skin depth is extremely small ( $\delta\sim 10^{-5}\text{cm}$ for $m_{a}=10^{-4}$eV ), restricting the current to a thin surface layer. As a result, the effective volume occupied by current flow is small, and the dissipated power from Joule heating is very low to be difficult to detect.
On the other hand, for a cylinder with low conductivity the induced current flows the bulk, and the resulting dissipated power $P$ becomes large enough to be detectable at low temperatures ( $T\sim 100\text{mK}$ ). We show that
$P\simeq 2.8\times 10^{-27}\mbox{W}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{m_{a}}{10^{-3}\text{eV}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (1)
with $\sigma=10m_{a}=10^{-2}$eV, where length L=100L=100cm and radius R=2R=2cm of the cylinder with electric permittivity $\epsilon=10$. The power $P$ depends slightly on $\sigma$ so that even with $\sigma=5\times 10^{-3}$eV, the power is still large $P\simeq 2.2\times 10^{-27}W$. Uniquely, the axion–electromagnetic coupling drives an oscillating current that permeates the entire cylindrical sample, in contrast to ordinary electromagnetic induction where currents are confined to the surface at high frequencies.
By adopting an appropriate conductivity, e.g. $\sigma\sim 5\times 10^{-3}\text{eV}$ at low temperature $\sim 100$mK and radius ( length ) of the cylinder 22cm ( 100100cm ), we show that using LC circuit tuned with a quality factor $Q=10^{6}$, dark matter axion can be probed in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$ with large signal-noise ratio $\sim 10$.
Our method using cylinder with low electric conductivity is also effective for so called, dark photon. The electromagnetic coupling with the dark photon gives rise to similar effect to the one in the axion that oscillating electric current is induced in the whole of cylinder.
@@
In this letter we show that, for a cylindrical sample in Fig.1 with low electrical conductivity, the Joule heating induced by axion dark matter can be large enough to be detectable at low temperatures $T\sim 100\text{mK}$. In this case, the dissipation occurs throughout the bulk of the cylinder, while for a good conductor with large conductivity, the dissipation is confined to a thin surface layer.
We focus on the axion mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$, using a cylinder of length $L=100\text{cm}$, radius $R\geq 10/m_{a}$, and conductivity $\sigma=10m_{a}=10^{-3}\text{-}10^{-2}\text{eV}$. In particular, we fix the radius R=2R=2cm thought out the paper as a practical value. As we show soon later, the condition of $\sigma m_{a}=\epsilon$ maximize the power dissipated inside the cylinder with electric permittivity $\epsilon$. Such a matter with low conductivity may be realized at low temperatures $\sim 100\text{mK}$, for instance by using semiconductors with suitable impurity doping.
@@
We calculate the axion-induced electric field to obtain oscillation electric current and its dissipation power $P$.
@@
where electric permittivity $\epsilon$ and electric current $\vec{J}$ are non vanishing inside the cylinder, while $\epsilon=1$ and $\vec{J}=0$ in vacuum. We have assumed trivial permeability $\mu=1$ of the cylinder.
@@
They oscillate as $E^{\prime}(B^{\prime})\propto\exp(-im_{a}t)$. It is easy to see that the first term of the electric field, $E^{\prime}(\rho)$, exists throughout the bulk of the cylinder, whereas the second term is confined to its surface; $J_{0}(bm_{a}\rho)/J_{0,1}(bm_{a}R)\propto\exp\big((R-\rho)m_{a}\text{Im}(b)\big)$ where $\text{Im}(O)$ denotes imaginary part of $O$.
Uniquely, the axion–electromagnetic coupling drives an oscillating electric field that permeates the entire cylindrical sample, in contrast to ordinary electromagnetic induction where the electric field is confined to the surface at high frequencies. A similar fact is also present in the case of dark photonkishimoto .
As shown later, the second term dominates the power dissipation in the limit of high conductivity $(\sigma\sim 10^{4}\text{eV})$, while the first term becomes dominant for low conductivity $(\sigma\sim 10^{-3}\text{eV})$.
@@
We remind that $y\equiv\sigma/m_{a}$ and $x\equiv m_{a}R$ and $b=b(y)$ is the function of $y$. $P$ is a complicated function in $x$ and $y$, or $m_{a}$ and $R$. But we show below that $P$ becomes a simple function in $x$ and $y$ when we take the limit of large conductivity, $\sigma\sim 10^{4}$eV or small $\sigma\sim 10^{-3}$eV.
It is easy to confirm that in the limit of $\sigma\to\infty$ ( $y\to\infty$ ), $P\propto m_{a}\rho_{a}V(B_{0}/m_{a}M)^{2}|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}$ with $M\equiv\pi f_{a}m_{a}/(g_{\gamma}\alpha)$ and the volume $V=2\pi RL\delta$, within which the electric current flows. $\delta=\sqrt{2/m_{a}\sigma}$ denotes skin depth. The current flows only in the surface with depth $\delta$ of the cylinder. It is coincide with the previous result kishimoto . In the limit, the main contribution comes from the second term in $U$. We note that $|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}\simeq 1$ for $x>5$.
The skin depth for general $\sigma$ is given by $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}=\Big(m_{a}(\epsilon^{2}+(\sigma/m_{a})^{2})^{1/4}\sin(\theta/2)\Big)^{-1}>\delta=\sqrt{2/m_{a}\sigma}$. In this paper, assuming typical value $\epsilon=10$ of semiconductor, we take low electric conductivity $\sigma\sim\epsilon m_{a}\sim 10m_{a}$ ( $y\sim 10$ ) at temperature $\sim 100$mK and consider the mass region $m_{a}=10^{-3}\mbox{-}10^{-4}\mbox{eV}$. The matter may be fabricated of for instance a semiconductor with appropriate doping of impurities.
In such a matter with low electric conductivity, we find that the main contribution to $P$ comes from the first term in $U$. Actually, the following quantities are monotonically increasing function in $x$,
@@
$P$ takes the maximal value at $y=\sigma/m_{a}=\epsilon$ and $P(\epsilon,y=\epsilon)\propto\epsilon^{-1}$. Thus, it is favorable to take the conductivity $\sigma=\epsilon m_{a}$. Because we do not know the value of the axion mass $m_{a}$, the value $\sigma=\epsilon m_{a}$ is unknown. But when we put $\sigma=\epsilon m_{a}$ in the formula, the dependence of $P$ on $m_{a}$ becomes simple; $P\propto m_{a}$. In the actual search of the axion mass in the range $10^{-4}\text{-}10^{-3}$eV, when we use semiconductor with $\epsilon\sim 10$, it is sufficient for the axion detection to take the value $\sigma\sim 5\times 10^{-3}$eV. Although it does not lead to the maximal power $P$, the predicted signal-noise ratios are large enough for the detection. Hereafter we take $\sigma=\epsilon m_{a}$ and $\epsilon=10$ for simplicity.
Numerically, we find
@@
with $\epsilon=10$, $y=10$ ( $\sigma=10m_{a}$ ) and $x\simeq 10$ ( R=2R=2cm ). Obviously, $P$ in eq.(20) is proportional to $R^{2}=x^{2}/m_{a}^{2}$, that is, the cross-sectional area $\pi R^{2}$ of the cylinder. The energy dissipation takes place in the bulk volume $\propto R^{2}L$, not merely in the surface, resulting in substantially enhanced dissipation power $P$.
@@
with $y=\sigma/m_{a}=10^{8}$ and $x=m_{a}R\simeq 10$. It is generated in the surface with the skin depth $\delta\sim 10^{-5}$cm of the cylinder. We find that the power $P(\sigma=10^{-3}\text{eV})$ with low conductivity, is three order of magnitude larger than $P_{h}(\sigma=10^{4}\text{eV})$ with high conductivity. That is, $P(\sigma=10^{-3}\text{eV})>10^{3}P(\sigma=10^{4}\text{eV})$. It is the reason why we use such a cylinder with low conductivity $\sigma\sim 10^{-3}$eV.
We compare $P$ with thermal noise $P_{t}=T\delta\omega/2\pi$ with the width $\delta\omega=10^{-6}m_{a}$ of the microwave frequency,
$P_{t}\simeq 3.3\times 10^{-20}W\Big(\frac{T}{100\rm mK}\Big)\Big(\frac{m_{a}}{10^{-4}\rm eV}\Big)$ (24)
We have
@@
$\frac{P(\sigma=10^{-3}\text{eV})}{P_{t}}\simeq 8.5\times 10^{-9}g_{\gamma}^{2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{10}{\epsilon}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (25)
We use LC circuit with resistance $R_{c}$, inductance $L_{c}$ and capacitance $C_{c}$ to detect the power $P$ induced by the axion. When we tune the circuit with resonant frequency $\omega=\sqrt{1/L_{c}C_{c}}=m_{a}$ and a quality factor $Q$ such as $Q=R_{c}^{-1}\sqrt{L_{c}/C_{c}}=10^{6}$, the signal to noise ( S/N ) ratio is given by
@@
$\frac{P(\sigma=10^{-3}\text{eV})}{P_{t}}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 10\,g_{\gamma}^{2}\Big(\frac{m_{a}}{10^{-4}\rm eV}\Big)^{1/2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\sqrt{\frac{t_{ob}}{60\rm s}}$ (26)
with $y=10$ and $x\simeq 10$, where $t_{ob}$ denotes the observation time.
When we change electric conductivity $\sigma=10^{-3}$eV slightly such as $\sigma=5\times 10^{-4}$eV, the S/N ratio $\simeq 8.2$. The dependence on the conductivity is very weak. But, when axion mass is smaller such as $m_{a}=10^{-5}$eV, it is favorable to use smaller $\sigma=10^{-4}$eV to maximize $P$. Then, it leads to
$\frac{P(\sigma=10^{-4}\text{eV})}{P_{t}}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 3.2g_{\gamma}^{2}\,\Big(\frac{m_{a}}{10^{-5}\rm eV}\Big)^{1/2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\sqrt{\frac{t_{ob}}{60\rm s}},$ (27)
with $y=10$ and $x\simeq 1$.
Furthermore, when axion mass is larger such as $m_{a}=10^{-3}$eV, it is favorable to use $\sigma=10^{-2}$eV to maximize $P$. It leads to
$\frac{P(\sigma=10^{-2}\text{eV})}{P_{t}}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 32\,g_{\gamma}^{2}\,\Big(\frac{m_{a}}{10^{-3}\rm eV}\Big)^{1/2}\Big(\frac{T}{100\rm mK}\Big)^{-1}\Big(\frac{L}{100\mbox{cm}}\Big)\Big(\frac{R}{2\mbox{cm}}\Big)^{2}\Big(\frac{B_{0}}{10\mbox{T}}\Big)^{2}\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)\sqrt{\frac{t_{ob}}{60\rm s}},$ (28)
with $y=10$ and $x\simeq 100$. Even when $\sigma=5\times 10^{-3}$eV, the S/N ratio is $\simeq 26$. We find that the dependence on the conductivity is also weak.
It turns out that we can search the region of the axion mass $m_{a}=10^{-4}\text{-}10^{-3}$eV with large S/N ratio. In particular, even by fixing $\sigma=5\times 10^{-3}$eV and R=2R=2cm, the S/N ratios are $\simeq 8.2$ for $m_{a}=10^{-4}$eV and $\simeq 26$ for $m_{a}=10^{-3}$eV. The dependence of S/N ratio on the conductivity $\sigma$ is weak so that even if the conductivity of the sample is not uniform in the large sample ( L=1L=1m and R=2R=2cm ), the signal of the axion dark matter could be found in the method.

### v2

@@ Abstract
A dark matter axion with mass $m_{a}$ induces an oscillating electric field in a cylindrical sample placed under a magnetic field $B$ parallel to the cylinder axis. When the cylinder is made of a highly electrically conductive material, the induced oscillating current flows only at the surface. In contrast, if the cylinder is composed of a material with small conductivity, e.g. $\sigma=10^{-3}\text{eV}$, the electric current flows inside the bulk of the cylinder. Within the QCD axion model, the current $I$ is estimated as $I(m_{a}=10^{-4}\text{eV})\simeq 1.9\times 10^{-14}\text{A}g_{\gamma}(R/6\text{cm})^{2}(\sigma/10^{-3}\text{eV})(B/10\text{T})(10/\epsilon)(\rho_{a}/0.3\text{GeVcm}^{-3})$ with radius $R$, length $L$, permittivity $\epsilon=10$ of the cylinder and axion energy density $\rho_{a}$, where $g_{\gamma}$ is model dependent parameter; $g_{\gamma}(\text{KSVZ})=-0.96$ and $g_{\gamma}(\text{DFSZ})=0.37$. Using an LC circuit with a quality factor $Q=10$, the signal-to-noise ratio at temperature T=100T=100mK and observation time $t_{ob}=60$second is $\sim 0.5(Q/10)\sqrt{(m_{a}/10^{-4}\text{eV})(100\text{mK}/T)(t_{ob}/60\text{s})}$. The detection of dark matter axions is feasible in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$.
@@
In this letter we propose a new method for axion detection using a cylindrical sample with small electrical conductivity, $\sigma=10^{-3}\text{-}10^{-2}\text{eV}$ at low temperature 100100mK. Our target is the mass range $m_{a}=10^{-4}\mbox{-}10^{-3}$eV, which is difficult to detect with resonant cavity.
@@
In general, such currents are confined within the skin depth $\delta$. For highly conductive materials, e.g. $\sigma=10^{4}\text{eV}$, the skin depth is extremely small ( $\delta\sim 10^{-5}\text{cm}$ for $m_{a}=10^{-4}$eV ), restricting the current to a thin surface layer due to the induction effect. Additionally, electric field induced in the conducting material is suppressed by the factor $\sqrt{m_{a}/\sigma}\sim 10^{-4}$, compared with the one induced in vacuum. As a result, the electric current is not large enough to be detectable.
On the other hand, the two suppression factors do not arise for electromagnetic fields interacting with axion, Especially, for the cylinder sample with small conductivity, the induced current mainly flows the bulk, and the resulting electric current $I$ becomes large enough to be detectable at low temperatures ( $T\sim 100\text{mK}$ ). We show that when the axion mass $m_{a}=10^{-4}$eV,
$I\simeq 2.1\times 10^{-15}\text{A}g_{\gamma}\Big(\frac{R}{2\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{10\text{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (1)
with $\sigma=10m_{a}=10^{-3}$eV, where $R$ denotes the radius of the cylinder with permittivity $\epsilon=10$. The current is maximized by choosing electrical conductivity $\sigma=\epsilon m_{a}$, for instance, $\sigma=10^{-3}\text{eV}(m_{a}/10^{-4}\text{eV})$ when $\epsilon=10$. $I$ is proportional to the surface area $R^{2}$. The oscillating current flows the entire cylindrical sample, in contrast to ordinary electromagnetic induction where currents with high frequencies are confined to the surface.
In the actual measurement we compare the current $I$ with current $I_{n}$ of thermal noise. When the noise current $I_{n}$ is smaller, the real signal current $I$ can be observed more easily. Thus, it is favorable that the noise $I_{n}$ is small. We adopt cylinder sample with small conductivity, e.g. $\sigma=10^{-3}$eV, which causes large resistance $R_{c}$ of the cylinder. Because $I_{n}\propto 1/\sqrt{R_{c}}$, it is expected that the thermal noise $I_{n}$ is small for large $R_{c}$. It is an advantage of using the sample with small conductivity $\sigma=10^{-3}$eV for the axion detection. As we show later, we have
$\frac{I}{I_{n}}\simeq 0.43\times 10^{-4}g_{\gamma}\sqrt{\frac{100\rm mK}{T}}\sqrt{\frac{\sigma}{10^{-3}\text{eV}}}\sqrt{\frac{10^{-4}\text{eV}}{m_{a}}}\sqrt{\frac{L}{100\mbox{cm}}}\Big(\frac{R}{2\mbox{cm}}\Big)\Big(\frac{B_{0}}{10\mbox{T}}\Big)\left(\frac{10}{\epsilon}\right)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (2)
with length $L$ of the cylinder.
By adopting an appropriate conductivity, e.g. $\sigma\sim 5\times 10^{-3}\text{eV}$ at low temperature $\sim 20$mK and radius ( length ) of the cylinder 66cm ( 100100cm ), we show that using parallel LC circuit tuned with a quality factor $Q=10$, dark matter axion can be probed in the mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$ with signal-noise ratio of the electric current bigger than $10$.
Our method using cylinder with small electrical conductivity is also effective for so called, dark photon. The electromagnetic coupling with the dark photon gives rise to similar effect to the one in the QCD axion that oscillating electric current is induced in the whole of cylinder.
@@
In this letter we show that, for a cylindrical sample in Fig.1 with small electrical conductivity, the electric current induced by axion dark matter can be large enough to be detectable at low temperatures $T\sim 100\text{mK}$. In this case, the current flows throughout the bulk of the cylinder, while for a good conductor with large conductivity, the current is confined to a thin surface layer.
We focus on the axion mass range $m_{a}=10^{-4}\text{-}10^{-3}\text{eV}$, using a cylinder of length $L=100\text{cm}$, radius $R\geq 2$ cm, and conductivity $\sigma=10m_{a}=10^{-3}\text{-}10^{-2}\text{eV}$. As we show soon later, the condition of $\sigma/m_{a}=\epsilon$ maximize the electric field induced in the cylinder with permittivity $\epsilon$. Such a matter with small conductivity may be realized at low temperatures $\sim 100\text{mK}$, for instance by using semiconductors with impurity doping appropriately.
@@
We calculate the axion-induced electric field $E^{\prime}$ to obtain oscillating electric current.
@@
where permittivity $\epsilon$ and electric current $\vec{J}$ are non vanishing inside the cylinder, while $\epsilon=1$ and $\vec{J}=0$ in vacuum. We have assumed trivial permeability $\mu=1$ of the cylinder.
@@
They oscillate as $E^{\prime}(B^{\prime})\propto\exp(-im_{a}t)$. Obviously, the first term of the electric field, $E^{\prime}(\rho)$, exists throughout the bulk of the cylinder, whereas the second term is confined to its surface; $J_{0}(bm_{a}\rho)/J_{0,1}(bm_{a}R)\propto\exp\big((R-\rho)m_{a}\text{Im}(b)\big)$ as $\rho\to\infty$ where $\text{Im}(O)$ denotes imaginary part of $O$.
Uniquely, the axion–electromagnetic coupling drives an oscillating electric field that permeates the entire cylindrical sample, in contrast to ordinary electromagnetic induction, where the high frequency electric field is confined to the surface. A similar fact is also present in the case of dark photonkishimoto .
As shown later, the second term dominates the electric power ( Joule heat ) in the limit of large conductivity $(\sigma\sim 10^{4}\text{eV})$, while the first term becomes dominant for small conductivity $(\sigma\sim 10^{-3}\text{eV})$.
@@
We remind that $y\equiv\sigma/m_{a}$ and $x\equiv m_{a}R$ and $b=b(y)$ is the function of $y$. $P$ is a complicated function in $x$ and $y$, or $m_{a}$ and $R$. But we show below that $P$ becomes a simple function in $x$ and $y$ when we take the limit of large conductivity, $\sigma\sim 10^{4}$eV or small $\sigma\sim 10^{-3}$eV. We note that electric field is written such that $E^{\prime}(\rho)=E_{a}U$.
It is easy to confirm that in the limit of $\sigma\to\infty$ ( $y\to\infty$ ), $P\propto m_{a}\rho_{a}V(B_{0}/m_{a}M)^{2}|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}$ with $M\equiv\pi f_{a}/(g_{\gamma}\alpha)$ and the volume $V=2\pi RL\delta$, within which the electric current flows. $\delta=\sqrt{2/m_{a}\sigma}$ denotes skin depth. The current flows only in the surface with the depth $\delta$ of the cylinder. It is coincide with the previous result kishimoto . In the limit, the main contribution comes from the second term in $U$. We note that $|H_{1}^{(1)}(x)/H_{0}^{(1)}(x)|^{2}\simeq 1$ for $x>5$.
The skin depth for general $\sigma$ represented by the second term of $U(x,y,z)$ in eq(18) is given by $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}=\Big(m_{a}(\epsilon^{2}+(\sigma/m_{a})^{2})^{1/4}\sin(\theta/2)\Big)^{-1}>\delta=\sqrt{2/m_{a}\sigma}$. In this paper, assuming typical value $\epsilon=10$ in semiconductor, we take small electric conductivity $\sigma\sim\epsilon m_{a}\sim 10m_{a}$ ( $y\sim 10$ ) and consider the mass region $m_{a}=10^{-3}\mbox{-}10^{-4}\mbox{eV}$. Then, the skin depth $\big(m_{a}\mbox{Im}(b(y))\big)^{-1}\simeq 0.1\text{cm}(10^{-4}\text{eV}/m_{a})$, smaller than the radius R=2R=2cm examined in our paper. On the other hand, the first term of $U$ is present in the whole of the cylinder, not confined at the surface.
In such a cylinder with small electrical conductivity, we find that the main contribution to $P$ comes from the first term in $U$. Actually, the following quantities are monotonically increasing function in $x$,
@@
$P$ takes the maximal value at $y=\sigma/m_{a}=\epsilon$ and $P(\epsilon,y=\epsilon)\propto\epsilon^{-1}$. Thus, it is favorable to take the conductivity $\sigma=\epsilon m_{a}$. Because we do not know the value of the axion mass $m_{a}$, the value $\sigma=\epsilon m_{a}$ is unknown. But when we put $\sigma=\epsilon m_{a}$ in the formula, the dependence of $P$ on $m_{a}$ becomes simple; $P\propto m_{a}$. In the actual search of the axion mass in the range $10^{-4}\text{-}10^{-3}$eV, it is sufficient for the axion detection to take the value $\sigma\sim 5\times 10^{-3}$eV with $\epsilon\sim 10$. Although it does not lead to the maximal power $P$, the predicted signal-noise ratios are large enough for the detection. Hereafter we take $\sigma=\epsilon m_{a}$ and $\epsilon=10$ for simplicity.
Numerically, noting $|E_{a}|=|g_{a\gamma\gamma}a(t)B_{0}|\simeq 1.6\times 10^{-36}\text{GeV}^{2}g_{\gamma}(B_{0}/10\text{T})$, we find
@@
with $\epsilon=10$, $y=10$ ( $\sigma=10m_{a}$ ) and $x\simeq 10$ ( R=2R=2cm ). Obviously, $P$ in eq.(21) is proportional to $R^{2}=x^{2}/m_{a}^{2}$, that is, the cross section $\pi R^{2}$ of the cylinder. The energy dissipation takes place in the bulk volume $\propto R^{2}L$, not merely in the surface, resulting in substantially enhanced dissipation power $P$.
@@
with $y=\sigma/m_{a}=10^{8}$ and $x=m_{a}R\simeq 10$. It is generated in the surface with the skin depth $\delta\sim 10^{-5}$cm of the cylinder. We find that the power $P(\sigma=10^{-3}\text{eV})$ with small conductivity, is three order of magnitude larger than $P(\sigma=10^{4}\text{eV})$ with large conductivity. That is, $P(\sigma=10^{-3}\text{eV})>10^{3}P(\sigma=10^{4}\text{eV})$. It is the reason why we use such a cylinder with small conductivity $\sigma\sim 10^{-3}$eV.
There are two reasons of the small power $P(\sigma=10^{4}\text{eV})$ in the cylinder. One is the suppression of electric field inside conductor with large conductivity $\sigma=10^{4}$eV, that is, $\sqrt{m_{a}/\sigma}|E_{a}|$ aiwazaki ; kishimoto compared with $|E_{a}|/2\epsilon$ in material with small conductivity $\sigma=10^{-3}$eV. The other one is the presence of the skin depth $\delta=\sqrt{2/m_{a}\sigma}$. The area $2\pi R\delta$ of the current flowing is much smaller than that $\pi R^{2}$ in material with low conductivity. Therefore, even if large conductivity $\sigma$ amplifies electric current by the factor $\sigma$, both the suppression of electric field by the factor $\sqrt{m_{a}/\sigma}$ and the area of current flow by the factor $\sqrt{1/m_{a}\sigma}$, cause the power $P(\sigma=10^{4}\text{eV})$ smaller than the power $P(\sigma=10^{-3}\text{eV})$. It is roughly,
$\frac{P(\sigma_{h}=10^{4}\text{eV})}{P(\sigma=10^{-3}\text{eV})}\sim\frac{R\delta\sigma_{h}(\sqrt{m_{a}/\sigma_{h}})^{2}}{R^{2}\sigma(1/\epsilon^{2})}=\frac{\delta m_{a}\epsilon^{2}}{R\sigma}\sim 10^{-3}$ (25)
with R=1R=1cm, $\epsilon=10$ and $m_{a}=10^{-4}$eV. It is coincident with the above calculations in eq(22) and eq(23).
The large power $P(\sigma=10^{-3}\text{eV})$ arises mainly from large electric current $I$ flowing in the bulk of the cylinder. The electric current $I$ is,
$I(\sigma=10^{-3}\text{eV})=\frac{\sigma|E_{a}|S}{2\epsilon}\simeq 2.1\times 10^{-15}\text{A}g_{\gamma}\Big(\frac{R}{2\text{cm}}\Big)^{2}\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{B_{0}}{10\text{T}}\Big),$ (26)
where the electric field $E=Re(E^{\prime})=|E_{a}|/2\epsilon=|E_{a}|/20$ is the one in the first term with $y=10$ in eq(14). We detect the axion by measuring the large electric current $I$.
The point causing the large current $I$ is that electric field $E$ inside of the cylinder with small conductivity, e.g. $\sigma=10^{-3}$eV has no suppression factor $\sqrt{m_{a}/\sigma}\simeq 10^{-4}$, which is present when the cylinder with large conductivity $\sigma=10^{4}$eV. Additionally, the current flows in the whole of the cylinder, not just in the surface. The facts make the large electric current $I$ to be observable.
There is another point which makes axion detection feasible. In the actual observation, we need to take into account thermal noise ( Johnson-Nyquist noise ). The electric current $I_{n}$ by the thermal noise is given by $I_{n}=\sqrt{2T\delta\omega/\pi R_{c}}$ with temperature T=100T=100mK and the width $\delta\omega$ of the frequency; $\delta\omega=10^{-6}m_{a}=10^{-10}(m_{a}/10^{-4}\text{eV})$eV. $R_{c}=L/(\sigma S)$ with $S=\pi R^{2}$ denotes electric resistance of the cylinder; $R_{c}=0.16(L/100\text{cm})(2\text{cm}/R)^{2}$. Then we find
$I_{n}=4.9\times 10^{-11}\text{A}\sqrt{\Big(\frac{T}{100\rm mK}\Big)\Big(\frac{m_{a}}{10^{-4}\rm eV}\Big)\Big(\frac{\sigma}{10^{-3}\text{eV}}\Big)\Big(\frac{100\text{cm}}{L}\Big)\Big(\frac{R}{2\text{cm}}\Big)^{2}}.$ (27)
Because the noise $I_{n}\propto 1/\sqrt{R_{c}}$, the noise is small for the large resistance $R_{c}\simeq 0.16$ of the cylinder. It should be stressed that the small noise $I_{n}$ is caused by small electrical conductivity $\sigma=10^{-3}$eV of the cylinder. Thus, we obtain the ratio of the current $I$ to the noise $I_{n}$,
@@
$\frac{I(\sigma=10^{-3}\text{eV})}{I_{n}(\sigma=10^{-3}\text{eV})}\simeq 0.43\times 10^{-4}g_{\gamma}\sqrt{\frac{100\rm mK}{T}}\sqrt{\frac{\sigma}{10^{-3}\text{eV}}}\sqrt{\frac{10^{-4}\text{eV}}{m_{a}}}\sqrt{\frac{L}{100\mbox{cm}}}\Big(\frac{R}{2\mbox{cm}}\Big)\Big(\frac{B_{0}}{10\mbox{T}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (28)
In order to amplify the above ratio $I/I_{n}$, we use, for instance, parallel LC circuit with resistance $R_{c}$, inductance $L_{c}$ and capacitance $C_{c}$. ( The resistance $R_{c}$ in the parallel LC circuit is almost equal to the one of the cylinder. ) When we tune the circuit with resonant frequency $\omega=\sqrt{1/L_{c}C_{c}}=m_{a}$ and a quality factor $Q$ such as $Q=R_{c}\sqrt{C_{c}/L_{c}}$, the above ratio is amplified such as $Q(I/I_{n})$. In actual experiment, it is difficult to have large quality factor $Q$ such as $Q>10^{3}$. So we assume $Q=10$ as a realistic value and measure the electric current in the coil of the LC circuit. Furthermore, in the measurement of the oscillating current, we take the frequency width $\delta\omega=10^{-6}m_{a}$ and observation time $60$ seconds. Then, the actual noise is reduced as $I_{n}/\sqrt{\delta\omega\delta t_{ob}/(2\pi)}\simeq I_{n}/(1.2\times 10^{3}\sqrt{\frac{m_{a}}{10^{-4}\text{eV}}}\sqrt{\frac{t_{ob}}{60s}})$.
It follows that the signal-noise ( SN ) ratio is
@@
$\frac{I(\sigma=10^{-3}\text{eV})}{I_{n}(\sigma=10^{-3}\text{eV})}Q\sqrt{\frac{\delta\omega t_{ob}}{2\pi}}\simeq 0.51\,g_{\gamma}\Big(\frac{Q}{10}\Big)\sqrt{\frac{100\rm mK}{T}}\sqrt{\frac{t_{ob}}{60\rm\,s}}\sqrt{\frac{\sigma}{10^{-3}\text{eV}}}\sqrt{\frac{L}{100\mbox{cm}}}\Big(\frac{R}{2\mbox{cm}}\Big)\Big(\frac{B_{0}}{10\mbox{T}}\Big)\Big(\frac{\rho_{a}}{0.3\rm GeVcm^{-3}}\Big)$ (29)
with $y=10$ and $x\simeq 10$, where $t_{ob}=60$ s denotes the observation time. Although the SN ratio does not explicitly involve the axion mass $m_{a}$, it implicitly depends on $m_{a}$ because $\sigma$ is taken as $\sigma=10m_{a}$. When we searches $m_{a}=10^{-3}$eV, it is favorable to have $\sigma=10^{-2}$eV. Then, the SN ratio is $\simeq 1.6$.
Therefore, it turns out that by the observation of the electric current $I$ with appropriate choices of, for instance, temperature T=20T=20mK and the radius R=6R=6cm, and observation time $t_{ob}=600$ s, SN ratio is larger than $10$. Without such a choice of the parameters, we can obtain large SN ratio as far as we tune quality factor $Q$ bigger than $10$.
We point out that when we search much larger axion mass $m_{a}>10^{-3}$eV, it is easier to detect such axions This is because electric current $I\propto\sigma$ and SN ratio $\propto\sqrt{\sigma}$ are much larger than those of $m_{a}=10^{-3}$eV with the appropriate choice of conductivity $\sigma=10m_{a}$.